OpenStudy (anonymous):

HELP PLEASE 1.Which equation represents the line passing through the points (2, −2) and (−5, 12)? 2x − y = 2 2x + y = 2 x + 2y = 2 x − 2y = 2 2.A conference room is in the shape of a rectangle. Its floor has a length of (x − 4) meters and a width of (3x − 1) meters. The expression below represents the area of the floor of the room in square meters. (x − 4)(3x − 1) Which of the following simplified expressions represents the area of the floor of the conference room in square meters? x2 − 13x + 4 3x2 − 13x + 4 3x2 − 11x + 4 x2 − 12x + 4

OpenStudy (anonymous):

slope can be found as \[m=\frac{12-(-2)}{-5-2}=\frac{14}{-7}=-2\] now using point (2,-2) use point slope form \[y-y _{0}=m(x-x _{0})\] \[y-(-2)=-2(x-2)\] simplify it and you will get your answer